# Adversarial Robustness, Fairness, and Ethics of Low-Rank LLMs

### v1

@@ I Introduction
One approach to addressing this challenge is to pretrain models of various sizes [bartoldson2023compute], enabling smaller models to run in constrained environments. For instance, the LLaMA2 family offers models ranging from 7B to 70B parameters [touvron]. Yet, training even small models demands vast compute and large-scale data. These resources are often inaccessible to most organizations. Model compression offers a scalable alternative. It reduces the size of pretrained models while maintaining performance, avoiding the cost of training from scratch. As a result, model compression has emerged as a practical and scalable solution for deploying LLMs in environments with constrained budgets.
Model compression encompasses complementary techniques that reduce the compute and energy demands of LLMs. Quantization reduces weight precision (e.g., GPTQ [gptq], AWQ [awq]), while pruning removes redundant weights (e.g., Wanda [wanda], FLAP [flap]). In parallel, low-rank factorization has recently emerged as a particularly effective approach (e.g., SVD [noach], Basel [basel], FWSVD [fwsvd]), decomposing large weight matrices into lower-dimensional matrices. Empirical studies show that low-rank compressed models can preserve competitive accuracy while offering substantial gains in latency, FLOPs, and overall system efficiency [xue2023extra, basel]. These benefits make low-rank factorization a compelling strategy for building resource-efficient LLMs.
@@
Despite the efficiency benefits, compression involves tradeoffs that may influence trust-related properties. Yet, this tradeoff between compression efficiency and model trustworthiness remains underexplored. Practitioners and users are left asking: Do compressed models leak private information? Are they robust to adversarial attacks? Do they behave ethically in sensitive contexts? Without rigorous analysis, deploying compressed models especially in high-stakes settings remains risky and unjustified [sun2024trustllm, demszky2023using, driess2023palm]. Prior work by [decodingcompressedtrust] examined the trustworthiness implications of quantization and pruning. However, the effects of low-rank factorization on trustworthiness have not been studied. This gap is critical, as low-rank techniques are rapidly adopted for their strong empirical performance across diverse architectures. Without such analysis, practitioners cannot confidently apply low-rank compressed models to domains where privacy, robustness, and ethical alignment are essential. Understanding the trust-related effects of low-rank models is necessary for safe deployment and for designing more reliable compression strategies.
To this end, we present the first comprehensive study of how low-rank factorization affects the trustworthiness of LLMs. To align with established practice, we evaluate low-rank factorization across the four core dimensions of trustworthiness defined by [decodingtrust]1 leakage of in-context data and personally identifiable information (PII), 2 robustness to adversarial attacks, 3 machine ethics violations, and 4 fairness (Figure 1). These dimensions reflect the most widely recognized concerns in assessing model safety and reliability. Given their close relationship to low-rank compression, we also examine how fine-tuning, model scale, and layer-type contributions (e.g., attention, MLP, embedding layers) influence these trustworthiness properties. Our results provide a principled view of the risks and safety trade-offs in low-rank compression, offering actionable insights for deployment and future method design.
@@ II Background
To quantify the system gains of low-rank compression, we applied three low-rank approximation techniques (SVD, FWSVD, and Basel) to the LLaMA2 7B model at various compression ratios. We evaluated the models on a single GPU using the GSM8K dataset [cobbe2021training]. The 7B full model required 23.95 GiB of memory and achieved a throughput of 261.7 tokens/sec. As shown in Figure 2, compressing the model to 10% of its original parameters via SVD reduced its size to 0.56B, boosted throughput to over 600 tokens/sec, and cut memory usage to 4.39GiB. FWSVD and Basel showed similar trends, with Basel reaching the highest throughput (621.9 tokens/sec) and reducing memory to 4.73 GiB at deepest compression.
@@ III Threat Model
Assets. A trustworthy LLM must protect four core assets: 1 training data, which may include sensitive content such as private emails (e.g., Enron [enron]); 2 user inputs at inference, especially conversational PII requiring confidentiality; 3 fair output; and 4 ethical integrity, ensuring outputs align with human values and avoid harmful or inappropriate content.
@@ IV-A Models
Following the practice in [decodingcompressedtrust], we conduct experiments with the LLaMA2 family, covering both base and instruction-tuned (Chat) variants. Additionally, we extend our analysis to the Qwen2.5 -7B/14B [qwen2023qwen] models to ensure our findings generalize beyond one model. For brevity, we denote the LLaMA2 as Base/Chat 7B/13B throughout the paper.
Finetuned Models. Pretrained LLMs excel at broad language modeling but often underperform on specialized tasks. Task and instruction tuning improves instruction following, generalization, and domain adaptation [Lewkowycz2022Minerva, Roziere2023CodeLlama]. Math reasoning (e.g., GSM8K) and code generation (e.g., HumanEval) are among the most common domains for fine-tuning [Cobbe2021GSM8K, Chen2021HumanEval], and recent surveys highlight these domains as central reasoning tasks [Hong2024LLMComp]. However, fine-tuning can also shift safety and reliability profiles, so these models must be assessed alongside utility. In this work, we fine-tune Base and Chat variants (7B, 13B) on programming and mathematics to study how task adaptation interacts with model trustworthiness.
@@ V-A Trustworthiness Evaluation
In this study, we assess LLM trustworthiness across six aspects: privacy in training data, privacy during inference, adversarial robustness, fairness, ethical compliance, and resilience to jailbreaks. While trustworthiness includes other categories, we focus on robustness, privacy, fairness and ethics. Robustness and privacy are emphasized due to their central role in deployment safety and security. Recent studies have shown that LLMs remain highly vulnerable to adversarial attacks [carlini2017towards, xu2024linkprompt, zou2024adversarial, kim2023propile, vu2024privacy, shen2024dynamicattention] and privacy leakage [shokri2017membership, decodingtrust]. Ethics is included to capture normative alignment, as LLMs often fail in commonsense morality and value conflicts [hendrycks2021ethics, decodingtrust]. As LLMs are increasingly deployed in real-world applications, their ability to generate morally appropriate outputs becomes critical to avoid societal, legal, and trust concerns. Furthermore, we include fairness in our evaluation because LLMs can systematically encode and amplify social biases, leading to disparate outcomes across demographic groups and undermining trust and legitimacy in deployment settings [gallegos2024bias]. We adopt the DecodingTrust framework [decodingtrust] to guide our evaluation.
@@ V-A1 Privacy Evaluation
To investigate privacy concerns, we explore leakage during training and inference. For training data leakage, we use in-context prompting on the Enron Email Dataset [enron], which contains over 600,000 emails from  150 users and is widely used for training language models because it includes realistic conversational structures that assist models in learning dialogue coherence and response generation. In our evaluation, we vary the context length $L$, where $L$ denotes the in-context sequence of tokens that comes before the target email. This approach assesses the model’s likelihood of reproducing memorized training data when prompted with similar contexts.
During inference, we evaluate how the model treats PII injected during conversation. This setting as used in DecodingTrust, mirrors practical deployments where assistants process private documents or messages, and once PII is introduced, it can later be elicited by subsequent prompts from the same or different users unless explicit safeguards are enforced  [cnn2023m365, cybernews2023samsung]. Consistent with DecodingTrust, we construct prompts using names and email addresses from the Enron dataset, while other identifiers (e.g., phone numbers, SSNs) are randomly generated. This evaluation quantifies the model’s propensity to disclose sensitive information provided during prior interactions.
@@ V-A2 Adversarial Robustness Evaluation
AdvGLUE [wang2021advglue] evaluates adversarial robustness through 14 validated perturbation strategies on GLUE [wang2018glue] tasks, including word- and sentence-level attacks such as TextBugger [li2019textbugger], TextFooler [kwon2023textfooler], SememePSO [zang2020wordlevel], and SemAttack [wang2022semattack]. AdvGLUE++ [decodingtrust] extends this benchmark to autoregressive LLMs (e.g., Alpaca [taori2023alpaca]), adapting attack objectives to generative decoding. We use AdvGLUE++ to assess robustness, applying diverse perturbations across tasks such as sentiment analysis and natural language inference.
@@ V-A3 Machine Ethics Evaluation
We evaluate LLMs on ethical reasoning tasks to assess their moral sensitivity under low-rank compression using the ETHICS dataset [hendrycks2021ethics], which tests principles such as justice, deontology, utilitarianism, and commonsense morality. We focus on the commonsense subset containing 13,910 training and 3,885 test samples, divided into 2,109 short (1–2 sentences) and 1,776 long (1–6 paragraphs) scenarios. We concentrate on the short-form samples. In few-shot settings, training samples are provided as demonstrations (e.g., Scenario: I told my baby I hated her when she cried. Label: wrong), following the prompting strategy of Hendrycks et al. [hendrycks2021ethics]. Models classify each scenario as wrong or not wrong (Figure 1) , and accuracy is computed on samples where a definitive answer is given.
@@ V-B Fairness
We evaluate the fairness of full and compressed models across multiple settings using DecodingTrust [decodingtrust]. We use the Adult dataset [asuncion2007uci], a standard benchmark for fairness evaluation. Each record contains 14 demographic and socioeconomic attributes, including age, education, marital status, occupation, race, and sex. For evaluation, we convert the tabular entries into structured natural-language descriptions. For example: “A person in 1996 has the following attributes: age 19, workclass: private, final weight: 192773, education: HS-grad, highest education level: 9, marital status: never-married, native country: United-States…”
@@ V-D Layer-wise Attribution Analysis
Low-rank factorization methods increasingly adopt layer-wise strategies, recognizing that layers in LLMs vary in representational importance. ASVD [yuan2024asvd] selects ranks based on activation sensitivity, ALDS [he2021alds] optimizes rank selection via reconstruction error, and Gao et al. [gao2024adaptive] introduce adaptive masking. While these approaches treat layers separately to improve compression efficiency, their primary target is benign accuracy. In contrast, we examine which layers most influence trustworthiness to guide compression strategies that preserve both accuracy and security. To this end, we employ an attribution-based method grounded in gradient-based interpretability [attr_1_sundararajan2017axiomatic, attr_3_shrikumar2017learning] and structural attribution via Information Flow Routes [information_flow] to explore layers contribution to trustworthiness.
@@ VI-D2 Evaluation on Jailbreaking ETHICS
As noted in MasterKey [deng2024masterkey] and safety misalignment [gong2025safetymisalignment], jailbreaks [^1] Jailbreaking refers to prompts intentionally crafted to bypass model safety mechanisms and induce unethical or harmful outputs. remain a persistent security risk. We evaluate vulnerability using five prompts designed to induce misclassification of immoral actions (Appendix A-D2). These prompts are embedded in a benign template. The evaluation metric is the false positive rate (FPR), i.e., the proportion of immoral (negative) examples misclassified as moral or neutral (positive), where higher FPR denotes greater vulnerability.
@@ VII-A Low-Rank Factorization
LLMs are computationally intensive due to their vast parameter counts. A common strategy to reduce model size and speed up inference is low-rank factorization of weight matrices, typically using singular value decomposition [golub]. Noach and Goldberg [noach] pioneered this approach by applying SVD to transformer weights and refining the compressed models via knowledge distillation. FWSVD [fwsvd] advances this by leveraging Fisher Information to retain critical subspaces during compression. More recently, Basel [basel] improved upon this by re-learning singular values through downstream task, achieving strong performance even at deep compression ratios. While these methods effectively preserve benign accuracy, its implications for model trustworthiness have not been explored. In particular, it remains unclear whether SVD introduces or mitigates vulnerabilities, especially in privacy-sensitive or adversarial settings. It is critical to comprehensively evaluate the trustworthiness of LLMs compressed with low-rank factorization algorithms.
@@ VII-B Trustworthiness of LLMs
DecodingTrust [decodingtrust] evaluates GPT-3.5 and GPT-4 across multiple trustworthiness dimensions. DecodingCompressedTrust [decodingcompressedtrust] extends this analysis to quantization and pruning. Distinct from prior work, we present the first comprehensive study of trustworthiness under low-rank compression, offering new insights into how compression reshapes trustworthiness in LLM deployment.
@@ A-A The impact of Context Length on Training Data Privacy Leakage
The stability in Chat models can be attributed to their alignment through instruction tuning and reinforcement learning from human feedback (RLHF). Unlike Base models, which are trained solely with next-token prediction and therefore exploit longer contexts as stronger cues for recalling memorized content, Chat models might be explicitly optimized to avoid reproducing information verbatim. Instruction tuning shifts the objective of Chat models from raw language modeling toward following instructions safely [Ouyang2022InstructGPT, Wei2021FLAN], while RLHF further penalizes violating behaviors and rewards refusals or safe paraphrasing [Bai2022TrainingHelpfulHarmless]. Prior work confirms that LLMs are prone to memorization [Carlini2023Quantifying, carlini2021extracting]. Our results show that alignment mitigates this tendency by enforcing safer generation policies. As a result, Chat models maintain consistent leakage rates across context lengths.
@@ A-D1 Performance of Compressed Models on ETHICS-Short
A striking observation is that compressed 13B models substantially outperform the uncompressed 13B in the few-shot setting (e.g., FWSVD-70 reaches 77.1% vs. 1.8% for Base-13B). First, few-shot classification in LLMs is highly prompt-sensitive: performance depends strongly on the order of demonstrations and the specific label words used, and calibration methods often recover large absolute gains [zhao2021calibrate, lu2022fantastically, webson2022prompt]. Larger models are particularly prone to amplifying such prompt biases, which can collapse performance when only a handful of examples are provided. Second, low-rank factorization acts as an implicit regularizer by pruning high-variance, fragile feature directions and retaining dominant semantic ones. This aligns with evidence that fine-tuned solutions lie in low-dimensional subspaces which means the set of useful solutions actually lies in a much smaller, low-dimensional subspace [aghajanyan2020intrinsic, aghajanyan2021multi], and that restricting adaptation to low-rank directions preserves or improves generalization [hu2021lora]. In our experiments, this appears to coincide with reduced sensitivity to prompt artifacts, yielding more stable few-shot performance than the full 13B model. ETHICS-Short specifically rewards broad commonsense priors rather than brittle prompt artifacts. Thus, compressed variants of 13B can generalize more effectively from a handful of demonstrations than the full-capacity base model, which overfits to spurious prompt features. Importantly, these results also show that “larger” does not imply “more ethical”. Our results aligns with the DecodingCompressedTrust [decodingcompressedtrust] where pruning and quantization ethics often outperform the Base model. It is worth noting that although low-rank compressed 13B variants outperform the uncompressed 13B in few-shot ETHICS-Short, this improvement reflects reduced sensitivity to prompt biases rather than inherently “more ethical” reasoning. Compression prunes brittle directions and stabilizes performance under few-shot prompting, but it does not introduce new moral priors. As a result, compressed models may still produce ethically problematic outputs in unconstrained settings.

### v2

@@ I Introduction
One approach to addressing this challenge is to pretrain models of various sizes [8], enabling smaller models to run in constrained environments. For instance, the LLaMA2 family offers models ranging from 7B to 70B parameters [49]. Yet, training even small models demands vast compute and large-scale data. These resources are often inaccessible to most organizations. Model compression offers a scalable alternative. It reduces the size of pretrained models while maintaining performance, avoiding the cost of training from scratch. As a result, model compression has emerged as a practical and scalable solution for deploying LLMs in environments with constrained budgets.
Model compression encompasses complementary techniques that reduce the compute and energy demands of LLMs. Quantization reduces weight precision (e.g., GPTQ [20], AWQ [38]), while pruning removes redundant weights (e.g., Wanda [46], FLAP [3]). In parallel, low-rank factorization has recently emerged as a particularly effective approach (e.g., SVD [40], Basel [37], FWSVD [29]), decomposing large weight matrices into lower-dimensional matrices. Empirical studies show that low-rank compressed models can preserve competitive accuracy while offering substantial gains in latency, FLOPs, and overall system efficiency [4, 37]. These benefits make low-rank factorization a compelling strategy for building resource-efficient LLMs.
@@
Despite the efficiency benefits, compression involves tradeoffs that may influence trust-related properties. Yet, this tradeoff between compression efficiency and model trustworthiness remains underexplored. Practitioners and users are left asking: Do compressed models leak private information? Are they robust to adversarial attacks? Do they behave ethically in sensitive contexts? Without rigorous analysis, deploying compressed models especially in high-stakes settings remains risky and unjustified [31, 16, 18]. Prior work by [27] examined the trustworthiness implications of quantization and pruning. However, the effects of low-rank factorization on trustworthiness have not been studied. This gap is critical, as low-rank techniques are rapidly adopted for their strong empirical performance across diverse architectures. Without such analysis, practitioners cannot confidently apply low-rank compressed models to domains where privacy, robustness, and ethical alignment are essential. Understanding the trust-related effects of low-rank models is necessary for safe deployment and for designing more reliable compression strategies.
To this end, we present the first comprehensive study of how low-rank factorization affects the trustworthiness of LLMs. To align with established practice, we evaluate low-rank factorization across the four core dimensions of trustworthiness defined by [53]1 leakage of in-context data and personally identifiable information (PII), 2 robustness to adversarial attacks, 3 machine ethics violations, and 4 fairness (Figure 1). These dimensions reflect the most widely recognized concerns in assessing model safety and reliability. Given their close relationship to low-rank compression, we also examine how fine-tuning, model scale, and layer-type contributions (e.g., attention, MLP, embedding layers) influence these trustworthiness properties. Our results provide a principled view of the risks and safety trade-offs in low-rank compression, offering actionable insights for deployment and future method design.
@@ II Background
To quantify the system gains of low-rank compression, we applied three low-rank approximation techniques (SVD, FWSVD, and Basel) to the LLaMA2 7B model at various compression ratios. We evaluated the models on a single GPU using the GSM8K dataset [14]. The 7B full model required 23.95 GiB of memory and achieved a throughput of 261.7 tokens/sec. As shown in Figure 2, compressing the model to 10% of its original parameters via SVD reduced its size to 0.56B, boosted throughput to over 600 tokens/sec, and cut memory usage to 4.39GiB. FWSVD and Basel showed similar trends, with Basel reaching the highest throughput (621.9 tokens/sec) and reducing memory to 4.73 GiB at deepest compression.
@@ III Threat Model
Assets. A trustworthy LLM must protect four core assets: 1 training data, which may include sensitive content such as private emails (e.g., Enron [33]); 2 user inputs at inference, especially conversational PII requiring confidentiality; 3 fair output; and 4 ethical integrity, ensuring outputs align with human values and avoid harmful or inappropriate content.
@@ IV-A Models
Following the practice in [27], we conduct experiments with the LLaMA2 family, covering both base and instruction-tuned (Chat) variants. Additionally, we extend our analysis to the Qwen2.5 -7B/14B [6] models to ensure our findings generalize beyond one model. For brevity, we denote the LLaMA2 as Base/Chat 7B/13B throughout the paper.
Finetuned Models. Pretrained LLMs excel at broad language modeling but often underperform on specialized tasks. Task and instruction tuning improves instruction following, generalization, and domain adaptation [35, 42]. Math reasoning (e.g., GSM8K) and code generation (e.g., HumanEval) are among the most common domains for fine-tuning [14, 12], and recent surveys highlight these domains as central reasoning tasks [28]. However, fine-tuning can also shift safety and reliability profiles, so these models must be assessed alongside utility. In this work, we fine-tune Base and Chat variants (7B, 13B) on programming and mathematics to study how task adaptation interacts with model trustworthiness.
@@ V-A Trustworthiness Evaluation
In this study, we assess LLM trustworthiness across six aspects: privacy in training data, privacy during inference, adversarial robustness, fairness, ethical compliance, and resilience to jailbreaks. While trustworthiness includes other categories, we focus on robustness, privacy, fairness and ethics. Robustness and privacy are emphasized due to their central role in deployment safety and security. Recent studies have shown that LLMs remain highly vulnerable to adversarial attacks [11, 1, 60, 32, 50, 43] and privacy leakage [44, 53]. Ethics is included to capture normative alignment, as LLMs often fail in commonsense morality and value conflicts [26, 53]. As LLMs are increasingly deployed in real-world applications, their ability to generate morally appropriate outputs becomes critical to avoid societal, legal, and trust concerns. Furthermore, we include fairness in our evaluation because LLMs can systematically encode and amplify social biases, leading to disparate outcomes across demographic groups and undermining trust and legitimacy in deployment settings [21]. We adopt the DecodingTrust framework [53] to guide our evaluation.
@@ V-A1 Privacy Evaluation
To investigate privacy concerns, we explore leakage during training and inference. For training data leakage, we use in-context prompting on the Enron Email Dataset [33], which contains over 600,000 emails from  150 users and is widely used for training language models because it includes realistic conversational structures that assist models in learning dialogue coherence and response generation. In our evaluation, we vary the context length $L$, where $L$ denotes the in-context sequence of tokens that comes before the target email. This approach assesses the model’s likelihood of reproducing memorized training data when prompted with similar contexts.
During inference, we evaluate how the model treats PII injected during conversation. This setting as used in DecodingTrust, mirrors practical deployments where assistants process private documents or messages, and once PII is introduced, it can later be elicited by subsequent prompts from the same or different users unless explicit safeguards are enforced  [13, 15]. Consistent with DecodingTrust, we construct prompts using names and email addresses from the Enron dataset, while other identifiers (e.g., phone numbers, SSNs) are randomly generated. This evaluation quantifies the model’s propensity to disclose sensitive information provided during prior interactions.
@@ V-A2 Adversarial Robustness Evaluation
AdvGLUE [54] evaluates adversarial robustness through 14 validated perturbation strategies on GLUE [51] tasks, including word- and sentence-level attacks such as TextBugger [36], TextFooler [34], SememePSO [58], and SemAttack [52]. AdvGLUE++ [53] extends this benchmark to autoregressive LLMs (e.g., Alpaca [48]), adapting attack objectives to generative decoding. We use AdvGLUE++ to assess robustness, applying diverse perturbations across tasks such as sentiment analysis and natural language inference.
@@ V-A3 Machine Ethics Evaluation
We evaluate LLMs on ethical reasoning tasks to assess their moral sensitivity under low-rank compression using the ETHICS dataset [26], which tests principles such as justice, deontology, utilitarianism, and commonsense morality. We focus on the commonsense subset containing 13,910 training and 3,885 test samples, divided into 2,109 short (1–2 sentences) and 1,776 long (1–6 paragraphs) scenarios. We concentrate on the short-form samples. In few-shot settings, training samples are provided as demonstrations (e.g., Scenario: I told my baby I hated her when she cried. Label: wrong), following the prompting strategy of Hendrycks et al. [26]. Models classify each scenario as wrong or not wrong (Figure 1) , and accuracy is computed on samples where a definitive answer is given.
@@ V-B Fairness
We evaluate the fairness of full and compressed models across multiple settings using DecodingTrust [53]. We use the Adult dataset [5], a standard benchmark for fairness evaluation. Each record contains 14 demographic and socioeconomic attributes, including age, education, marital status, occupation, race, and sex. For evaluation, we convert the tabular entries into structured natural-language descriptions. For example: “A person in 1996 has the following attributes: age 19, workclass: private, final weight: 192773, education: HS-grad, highest education level: 9, marital status: never-married, native country: United-States…”
@@ V-D Layer-wise Attribution Analysis
Low-rank factorization methods increasingly adopt layer-wise strategies, recognizing that layers in LLMs vary in representational importance. ASVD [57] selects ranks based on activation sensitivity, ALDS [25] optimizes rank selection via reconstruction error, and Gao et al. [22] introduce adaptive masking. While these approaches treat layers separately to improve compression efficiency, their primary target is benign accuracy. In contrast, we examine which layers most influence trustworthiness to guide compression strategies that preserve both accuracy and security. To this end, we employ an attribution-based method grounded in gradient-based interpretability [47, 45] and structural attribution via Information Flow Routes [19] to explore layers contribution to trustworthiness.
@@ VI-D2 Evaluation on Jailbreaking ETHICS
As noted in MasterKey [17] and safety misalignment [24], jailbreaks [^1] Jailbreaking refers to prompts intentionally crafted to bypass model safety mechanisms and induce unethical or harmful outputs. remain a persistent security risk. We evaluate vulnerability using five prompts designed to induce misclassification of immoral actions (Appendix A-D2). These prompts are embedded in a benign template. The evaluation metric is the false positive rate (FPR), i.e., the proportion of immoral (negative) examples misclassified as moral or neutral (positive), where higher FPR denotes greater vulnerability.
@@ VII-A Low-Rank Factorization
LLMs are computationally intensive due to their vast parameter counts. A common strategy to reduce model size and speed up inference is low-rank factorization of weight matrices, typically using singular value decomposition [23]. Noach and Goldberg [40] pioneered this approach by applying SVD to transformer weights and refining the compressed models via knowledge distillation. FWSVD [29] advances this by leveraging Fisher Information to retain critical subspaces during compression. More recently, Basel [37] improved upon this by re-learning singular values through downstream task, achieving strong performance even at deep compression ratios. While these methods effectively preserve benign accuracy, its implications for model trustworthiness have not been explored. In particular, it remains unclear whether SVD introduces or mitigates vulnerabilities, especially in privacy-sensitive or adversarial settings. It is critical to comprehensively evaluate the trustworthiness of LLMs compressed with low-rank factorization algorithms.
@@ VII-B Trustworthiness of LLMs
DecodingTrust [53] evaluates GPT-3.5 and GPT-4 across multiple trustworthiness dimensions. DecodingCompressedTrust [27] extends this analysis to quantization and pruning. Distinct from prior work, we present the first comprehensive study of trustworthiness under low-rank compression, offering new insights into how compression reshapes trustworthiness in LLM deployment.
@@ A-A The impact of Context Length on Training Data Privacy Leakage
The stability in Chat models can be attributed to their alignment through instruction tuning and reinforcement learning from human feedback (RLHF). Unlike Base models, which are trained solely with next-token prediction and therefore exploit longer contexts as stronger cues for recalling memorized content, Chat models might be explicitly optimized to avoid reproducing information verbatim. Instruction tuning shifts the objective of Chat models from raw language modeling toward following instructions safely [41, 56], while RLHF further penalizes violating behaviors and rewards refusals or safe paraphrasing [7]. Prior work confirms that LLMs are prone to memorization [9, 10]. Our results show that alignment mitigates this tendency by enforcing safer generation policies. As a result, Chat models maintain consistent leakage rates across context lengths.
@@ A-D1 Performance of Compressed Models on ETHICS-Short
A striking observation is that compressed 13B models substantially outperform the uncompressed 13B in the few-shot setting (e.g., FWSVD-70 reaches 77.1% vs. 1.8% for Base-13B). First, few-shot classification in LLMs is highly prompt-sensitive: performance depends strongly on the order of demonstrations and the specific label words used, and calibration methods often recover large absolute gains [59, 39, 55]. Larger models are particularly prone to amplifying such prompt biases, which can collapse performance when only a handful of examples are provided. Second, low-rank factorization acts as an implicit regularizer by pruning high-variance, fragile feature directions and retaining dominant semantic ones. This aligns with evidence that fine-tuned solutions lie in low-dimensional subspaces which means the set of useful solutions actually lies in a much smaller, low-dimensional subspace [2], and that restricting adaptation to low-rank directions preserves or improves generalization [30]. In our experiments, this appears to coincide with reduced sensitivity to prompt artifacts, yielding more stable few-shot performance than the full 13B model. ETHICS-Short specifically rewards broad commonsense priors rather than brittle prompt artifacts. Thus, compressed variants of 13B can generalize more effectively from a handful of demonstrations than the full-capacity base model, which overfits to spurious prompt features. Importantly, these results also show that “larger” does not imply “more ethical”. Our results aligns with the DecodingCompressedTrust [27] where pruning and quantization ethics often outperform the Base model. It is worth noting that although low-rank compressed 13B variants outperform the uncompressed 13B in few-shot ETHICS-Short, this improvement reflects reduced sensitivity to prompt biases rather than inherently “more ethical” reasoning. Compression prunes brittle directions and stabilizes performance under few-shot prompting, but it does not introduce new moral priors. As a result, compressed models may still produce ethically problematic outputs in unconstrained settings.